Use KVL to determine the Voltage Source Ez (V). E3 (V) = ? Is =2A E2 두 + V - - R4100 + 300 + E R2로L 8V 100 E3 두
Answer #1

E2 = 30v V. + RIA100 Ls= 2a E-Jov E3 R2원 vs =8v 빠 Apply K VL in mesh - E - E2 + Vrt Ve + E3 = 0 - 10-30 + 102 +8 + E3 = 0 ⇒